a steel plate measures '3 x 5'. If the pressure is 30lb/in2, calculate the force on the plate.
1
Expert's answer
2019-03-14T17:39:24-0400

Given: "x = 3~\text{ft};~y = 5~\text{ft};~p = 30~\frac{\text{lb}}{\text{in}^2}."

Since, by definition, pressure "p = \frac{F}{A}," where F is force and A is area, and area of a rectangular plate is "A = xy," solving for the force, we get:

"F = pA = pxy."

Entering the numerical values and converting the feet to inches (for canceling them out),

"F = 30~\frac{\text{lb}}{\text{in}^2}~*~(3*12)~\text{in}~*~(5*12)~\text{in} = 64800~lb,"

expressed in the pound-force units.